OpenStudy (anonymous):

A ball is thrown from a height of 123 feet with an initial downward velocity of 11/fts. The ball's height h (in feet) after t seconds is given by the following. h=123−11t-16t^2 How long after the ball is thrown does it hit the ground?

OpenStudy (anonymous):

@whpalmer4 Can you help me with this one?

OpenStudy (whpalmer4):

okay, here you need to find the values of \(t\) where \(h=0\)

OpenStudy (whpalmer4):

one of them will be negative, and can be disregarded

OpenStudy (anonymous):

So how would I do that?

OpenStudy (whpalmer4):

Same way you did the other one?

OpenStudy (whpalmer4):

\[0 = 123-11t-16t^2\]

OpenStudy (anonymous):

Would I be solving for h?

OpenStudy (anonymous):

I got t= -11/32 + 1/32sqrt7993 and t=-11/32 + -1/32sqrt7993

OpenStudy (whpalmer4):

no, you know h = 0, you're solving for values of t that make that true. and it looks like you got them correct :-)

OpenStudy (anonymous):

And the numerical values would be -3.34 and 2.45 ?

OpenStudy (anonymous):

-3.14**

OpenStudy (whpalmer4):

yes, that's right